# Systematic Search for FFPs in KMTNet Full-Frame Images. I. Photometry Pipeline

### v2

@@ 1 Introduction
Recently, Yang et al. 2024 (hereafter Y24) optimized the pySIS pipeline to be more automatic and more efficient. Therefore, together with more computational resources available, we initiated a project to develop a new KMTNet FFDI pipeline based on the Y24 pipeline and then utilize the produced photometric data to search for buried KMTNet FFP events. We name this project “Systematic Search for FFPs in KMTNet Full-Frame Images”. As the first paper of this series, we introduce the pipeline setups and the preliminary search results on a 1-year 1-deg2 subset of the full-frame images.
@@ 2 NEW FFDI Pipeline
Figure 1: An example BLG02 N-chip full-frame image taken by KMTC. The image has $9232\times 9216$ pixels. The red box includes a zoomed-in image of $\mathrm{470\times 470~{}pixel^{2}}$, which is the same as our stamp size for photometry in Section 2.2. The light green lines indicate the boundaries of the eight CCD read-out channels on the chip. Difference image algorithms are needed for such a dense field.
@@
KMTNet camera has (K, M, T, N) four chips, and each chip has 9232$\times$9216 pixels with an average pixel scale of $\sim 0.4$ arcsec. The field of view of each chip is about 1 $\mathrm{deg^{2}}$. Most images are taken in $I$ band, and 1/$1/1$ 11 of images are taken in $V$ band for source color measurements. KMTNet has $\sim 13~{}\rm{deg}^{2}$ of prime fields, (BLG01, BLG02, BLG03) and (BLG41, BLG42, BLG43), with $\sim 8^{\prime}$ shifts to fill the gaps between CCD chips (Kim et al., 2018a). We start by establishing our pipeline on a subset of the images. They are the $I$-band N-chip images of the BLG02 and BLG42 fields taken in 2018. The field has the highest event rate (see Figure 7 in Kim et al. 2018a). The N chip of BLG02 and BLG42 is centered at (R.A., Dec.) = (17:56:51.82, --29:35:02.28) and (17:56:34.72, --29:42:31.01), respectively. An example image is shown in Figure 1. The observing cadences of the BLG02 and BLG42 fields are $\Gamma=2~{}{\rm hr}^{-1}$ for KMTC02 and KMTC42 and $\Gamma=3~{}{\rm hr}^{-1}$ for KMTA02, KMTA42, KMTS02, and KMTS42333For simplification, we denote the KMTC-BLG02 field as KMTC02, and so on.. The subset has 14593 $I$-band images in total, including (2504, 2400, 2133, 2127, 2725, 2704) images from (KMTC02, KMTC42, KMTS02, KMTS42, KMTA02, KMTA42), respectively.
@@ 2.1.2 Coarse Transformations
Generally, finding the astrometric solution between two catalogs involves two steps: cross-matching stars and calculating transformations. The two steps depend on each other, meaning that at least one initial guess must be provided, and an iterative approach is necessary. We begin by estimating the initial transformations. The goal is to make the common stars in two catalogs get close enough to enable the cross-match.
For the initial transformation among KMTNet images from the same site-field combination, we simply apply pure translation. The reason is that the discrepancies between these images are primarily due to overall shifts in pointing. The translations are estimated by locating the peaks of the catalog-catalog star position cross-correlation functions. After applying the translation, the distances of common stars across different image catalogs are reduced to $\lesssim 2$ arcsec or $\lesssim 5$ pixels. This distance is sufficiently small compared to the average star distance of $\sim 18$ pixels within a bright star catalog.
Transforming the master KMTNet image frames to celestial coordinates requires more than a simple translation, as the stellar sphere and the rotation and optical distortion cannot be ignored. Fortunately, the translation approximation remains adequate on smaller scales. Therefore, we first divide the 1 deg2 field into $5\times 5$ sub-fields and find the translations separately. Next, we combine these translations using a 5th order Legendre polynomial function to obtain a global initial transformation for the entire 1 deg2 field. After the transformation, the distances of common stars between the master KMTNet catalogs and the Gaia catalog are $\lesssim 1$ arcsec.
@@ 2.1.3 Refined Transformations
After obtaining the initial transformations, the next step is to find common star pairs in the two catalogs, a process known as cross-matching. When conducting the cross-match between two catalogs, we search for pairs of stars that are mutual nearest neighbors and regard them as the same star. Star pairs with distances greater than $1\mathrm{arcsec}$ are excluded. The process is speeded up by the k-d tree algorithm using the scipy.spatial.KDTree package.
After identifying the common stars in two catalogs, we fit the transformation between their positions using Legendre polynomials. We then update the cross-match iteratively with the new transformation. The fit-and-match iteration converges in approximately 15 iterations. During the iteration, we gradually decrease the maximum allowed distance of the cross-match from 5 to 0.2 pixels (i.e., from $2^{\prime\prime}$ to $0.08^{\prime\prime}$). Moreover, the order of the Legendre polynomials is gradually increased until it reaches the maximum value of 20 to fit the complex distortion pattern.
@@ 2.2 Difference Image Analysis
We divide the full-frame images into smaller sub-image stamps to perform difference imaging and photometry. The stamp size should be small enough to prevent significant PSF and background changes but large enough to avoid poor PSF and poor convolution kernel constructions due to insufficient star numbers. Additionally, because the 8 horizontally arranged 1154-pixel width readout channels (see Figure 1) of each CCD chip have slightly different gains and flux offsets, we require the stamps to be mostly located in the same channel. By statistically analyzing the telescope’s pointing-induced horizontal shifts across all images (typically $<20$ pixels), we use the most frequent channel boundary positions as the reference for the stamp divisions to maximize the channel consistency.
@@ 4.1 New Microlensing Events
Table 2: PSPL Parameters for the Three New PSPL Microlensing Events
$\chi^{2}$/dof 11366/11366 9795/9795 11217/11217
$u_{0}$ $2.428\pm 0.049$ $0.136\pm 0.072$ $2.062\pm 0.014$
$t_{\rm E}$ (days) $2.62\pm 0.11$ $4.7\pm 2.1$ $14.31\pm 0.23$
$I_{\rm S,KMTC02}$ $16.6826\pm 0.0001$ $21.5\pm 0.6$ $17.0928\pm 0.0003$

### v3

@@ 1 Introduction
Recently, Yang et al. 2024 (hereafter Y24) optimized the pySIS pipeline to be more automatic and more efficient. Therefore, together with more computational resources available, we initiated a project to develop a new KMTNet FFDI pipeline based on the Y24 pipeline and then utilize the produced photometric data to search for buried KMTNet FFP events. We name this project “Systematic Search for FFPs in KMTNet Full-Frame Images.” As the first paper of this series, we introduce the pipeline setups and the preliminary search results on a 1-year 1-deg2 subset of the full-frame images.
@@ 2 NEW FFDI Pipeline
Figure 1: An example BLG02 N-chip full-frame image taken by KMTC. The image has $9216\times 9232$ pixels. The red box includes a zoomed-in image of $\mathrm{470\times 470~{}pixel^{2}}$, which is the same as our stamp size for photometry in Section 2.2. The light green lines indicate the boundaries of the eight CCD read-out channels on the chip. Difference image algorithms are needed for such a dense field.
@@
KMTNet camera has (K, M, T, N) four chips, and each chip has 9216$\times$9232 pixels with an average pixel scale of $\sim 0.4$ arcsec. The field of view of each chip is about 1 $\mathrm{deg^{2}}$. Most images are taken in $I$ band, and 1/$1/1$ 11 of images are taken in $V$ band for source color measurements. KMTNet has $\sim 13~{}\rm{deg}^{2}$ of prime fields, (BLG01, BLG02, BLG03) and (BLG41, BLG42, BLG43), with $\sim 8^{\prime}$ shifts to fill the gaps between CCD chips (Kim et al., 2018a). We start by establishing our pipeline on a subset of the images. They are the $I$-band N-chip images of the BLG02 and BLG42 fields taken in 2018. The field has the highest event rate (see Figure 7 in Kim et al. 2018a). The N chip of BLG02 and BLG42 is centered at (R.A., Dec.) = (17:56:51.82, --29:35:02.28) and (17:56:34.72, --29:42:31.01), respectively. An example image is shown in Figure 1. The observing cadences of the BLG02 and BLG42 fields are $\Gamma=2~{}{\rm hr}^{-1}$ for KMTC02 and KMTC42 and $\Gamma=3~{}{\rm hr}^{-1}$ for KMTA02, KMTA42, KMTS02, and KMTS42333For simplification, we denote the KMTC-BLG02 field as KMTC02, and so on.. The subset has 14593 $I$-band images in total, including (2504, 2400, 2133, 2127, 2725, 2704) images from (KMTC02, KMTC42, KMTS02, KMTS42, KMTA02, KMTA42), respectively.
@@ 2.1.2 Coarse Transformations
Generally, finding the astrometric solution between two catalogs involves two steps: cross-matching stars and calculating transformations. The two steps depend on each other, meaning that at least one initial guess must be provided, and an iterative approach is necessary. We begin by estimating the initial transformations. The goal is to make the mutual stars in two catalogs get close enough to enable the cross-match.
For the initial transformation among KMTNet images from the same site-field combination, we simply apply pure translation. The reason is that the discrepancies between these images are primarily due to overall shifts in pointing. The translations are estimated by locating the peaks of the catalog-catalog star position cross-correlation functions. After applying the translation, the distances of mutual stars across different image catalogs are reduced to $\lesssim 2$ arcsec or $\lesssim 5$ pixels. This distance is sufficiently small compared to the average star distance of $\sim 18$ pixels within a bright star catalog.
Transforming the master KMTNet image frames to celestial coordinates requires more than a simple translation, as the stellar sphere and the rotation and optical distortion cannot be ignored. Fortunately, the translation approximation remains adequate on smaller scales. Therefore, we first divide the 1 deg2 field into $5\times 5$ sub-fields and find the translations separately. Next, we combine these translations using a 5th order Legendre polynomial function to obtain a global initial transformation for the entire 1 deg2 field. After the transformation, the distances of mutual stars between the master KMTNet catalogs and the Gaia catalog are $\lesssim 1$ arcsec.
@@ 2.1.3 Refined Transformations
After obtaining the initial transformations, the next step is to find mutual star pairs in the two catalogs, a process known as cross-matching. When conducting the cross-match between two catalogs, we search for pairs of stars that are mutual nearest neighbors and regard them as the same star. Star pairs with distances greater than $1\mathrm{arcsec}$ are excluded. The process is speeded up by the k-d tree algorithm using the scipy.spatial.KDTree package.
After identifying the mutual stars in two catalogs, we fit the transformation between their positions using Legendre polynomials. We then update the cross-match iteratively with the new transformation. The fit-and-match iteration converges in approximately 15 iterations. During the iteration, we gradually decrease the maximum allowed distance of the cross-match from 5 to 0.2 pixels (i.e., from $2^{\prime\prime}$ to $0.08^{\prime\prime}$). Moreover, the order of the Legendre polynomials is gradually increased until it reaches the maximum value of 20 to fit the complex distortion pattern.
@@ 2.2 Difference Image Analysis
We divide the full-frame images into smaller sub-image stamps to perform difference imaging and photometry. The stamp size should be small enough to prevent significant PSF and background changes but large enough to avoid poor PSF and poor convolution kernel constructions due to insufficient star numbers. Additionally, because the 8 horizontally arranged 1152-pixel width readout channels (see Figure 1) of each CCD chip have slightly different gains and flux offsets, we require the stamps to be mostly located in the same channel. By statistically analyzing the telescope’s pointing-induced horizontal shifts across all images (typically $<20$ pixels), we use the most frequent channel boundary positions as the reference for the stamp divisions to maximize the channel consistency.
@@ 5.2 Prospect of an FFP Search on a Larger Scope and New AlertFinder System
Table 3: PSPL Parameters for the Three New PSPL Microlensing Events
$\chi^{2}$/dof 11366/11366 9795/9795 11217/11217
$u_{0}$ $2.428\pm 0.049$ $0.136\pm 0.072$ $2.062\pm 0.014$
$t_{\rm E}$ (days) $2.62\pm 0.11$ $4.7\pm 2.1$ $14.31\pm 0.23$
$I_{\rm S,KMTC02}$ $16.6826\pm 0.0001$ $21.5\pm 0.6$ $17.0928\pm 0.0003$